OpenStudy (anonymous):

can someone show me the steps to solving this equation –5 = s – 3

OpenStudy (anonymous):

Move the 3 to the other side so you get -5 + 3 = s -3 + 3 That then is -2 = s

OpenStudy (aroub):

you add 3 by both sides , -5+3=s+3-3, -2=s

OpenStudy (aroub):

TO both sides not BY both sides lol sorryyy!!